OpenStudy (x__hazel__x):

Determine the sum of the given probabilities. Are the events complementary? @kayders1997 P(red) = 2/5 ; P(blue) = 3/5 A. 1/2 The events are complementary. B. 1/2 The events are not complementary. C. 1 The events are complementary. D. 1 The events are not complementary.

OpenStudy (kayders1997):

First what is the sum when you add these together?

OpenStudy (x__hazel__x):

Umm.. \[\frac{ 2 }{ 5 } + \frac{ 3 }{ 5 } = \frac{ 5 }{ 5 } \]

OpenStudy (x__hazel__x):

Which is 1 :3

OpenStudy (x__hazel__x):

So im guessing A and B are crossed out!

OpenStudy (kayders1997):

Yes, now you need to think about is it complimentary which is like saying do you have only two outcomes, like when flippling a coin you can only gets heads or tails is that kinda the case here?

OpenStudy (x__hazel__x):

No, because you will get a better chance of pulling out blue!

OpenStudy (x__hazel__x):

D Right?

OpenStudy (kayders1997):

Yeah, I'm pretty sure thats right, at first I thought it was yes because there were only 2 colors but there is more of one than the other

OpenStudy (x__hazel__x):

Okay! Thank you!

OpenStudy (kayders1997):

No problem :)

OpenStudy (kayders1997):

Any others? :)